OpenStudy (anonymous):

medal and fan awarded solve the proportion x over x^2 - 2 = 2 over 3x + 5. What is the value of x? x = 2 x = -2 and x = 2 x = -1 and x = -4 x = 4

OpenStudy (jdoe0001):

ahemm, do you know how to factor? I assume so, otherwise it wouldn't make sense to give you this question

OpenStudy (anonymous):

Well I obviously need help either way or else I wouldn't be here.

OpenStudy (anonymous):

nvm I got help form elsewhere its x = −1 and x = −4

OpenStudy (jdoe0001):

yes, but I'm assuming you do know how to factor, I mean, if you haven't covered it, you should first, thus to get rid of the denominators multiply both sides like so $$ \cfrac{x}{\color{red}{x^2-2}}=\cfrac{2}{\color{blue}{3x+5}}\\ \cfrac{\color{blue}{3x+5}}{\color{red}{x^2-2}} \times \cfrac{x}{\color{red}{x^2-2}}=\cfrac{2}{\color{blue}{3x+5}} \times \cfrac{\color{blue}{3x+5}}{\color{red}{x^2-2}} $$ what would that give you?